OpenStudy (anonymous):

find the least possible value of alpha for which 1) sin (alpha-theta) =cos (alpha + theta) 2) sin ( theta + 2*alpha) = cos (alpha - theta )

ganeshie8 (ganeshie8):

they look tricky, what have you tried so far ?

OpenStudy (anonymous):

Hello @ganeshie8 , thanks for trying at least to answer. well .. I tried to draw both trigonmetric eqn ( cos x and sin x ) and assumed that theta to be always zero .. but im wondering if i shud refer to the trig ratios

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Well the least value for witch sin=cos is pi/4 If we put alpha+theta=alpha-theta => theta=0 & alpha=pi/4 I'm not sure about this though, maybe try something along the lines of sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B and see where it gets you. Sorry couldn't be much help.